A man has 7 nickels and 5 quarters in his pocket. He removes two coins without replacement. Find the probability that both are quarters.
Answer by stanbon(75887)

A man has 7 nickels and 5 quarters in his pocket. He removes two coins without replacement. Find the probability that both are quarters.
-----
# of ways to succeed: 5C2 = 10
# of possible outcomes: 12C2 = 66
====================================
P(2 quarters) = 10/66 = 5/33
=====
Cheers,
Stan H.
==================